More than 90% of people surveyed say they always wash their hands after using the public restrooms. But observers posted in public restrooms found that only 4679 of 6076 observed individuals actually washed their hands.

A.) What population does inference concern here?

B.) Explain clearly what the population proportion p represents in this setting.

C.) What is the numerical value of the sample proportion ^

p?
Answer #1

A) The number of people say that they always wash their hands after using the public restrooms.

B) The population proportion is 90%=0.90

C) The sample proportion is ,

p=X/n=4679/6076=0.7701